OpenStudy (anonymous):

Please help!! (3-x) = 2(x) x = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\(\Large 3-x=2x\) \(\Large \Rightarrow 2x+x=3\) \(\Large \Rightarrow 3x=3\) \(\Large \Rightarrow x=\frac{3}{3}=1\)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

ur welcome ^^